# AesFA: An Aesthetic Feature-Aware Arbitrary Neural Style Transfer

### 10 Dec 2023 v1

@@ 1 Introduction
Figure 1: Top: The Starry Night by Vincent Van Gogh. The shortcomings of the existing method: 1) Styles are not agnostic to spatial information, and 2) for ultra-high resolution style transfer, the effectiveness of the encoding networks is deficient, creating undesirable artifacts or suboptimal stylizations. Bottom: Compared with other NST methods, our method can faithfully transfer styles while ensuring the spatial information.
@@
2 Related Work
NST emerged with Gatys et al. (Gatys, Ecker, and Bethge 2016), but its optimization is computationally intensive. To deal with this issue, Johnson et al. (Johnson, Alahi, and Fei-Fei 2016) introduced perceptual losses for real-time processing. Subsequent work (Gatys et al. 2017; Ghiasi et al. 2017; Chen and Schmidt 2016; Ulyanov, Vedaldi, and Lempitsky 2017; Dumoulin, Shlens, and Kudlur 2016; Ulyanov et al. 2016) improved NST without sacrificing speed. However, all were limited to specific styles. Huang et al. (Huang and Belongie 2017) proposed Adaptive Instance Normalization (AdaIN) for arbitrary style transfer, which has been extended (Sheng et al. 2018; Kotovenko et al. 2019; Jing et al. 2020; Shen, Yan, and Zeng 2018; Li et al. 2017) for successful style transfer onto content images. Chandran et al. (Chandran et al. 2021) improved AdaIN with AdaConv for structure-aware style transfer. AdaConv simultaneously adapts statistical and structural styles. However, AdaConv’s convolution kernels and biases incur high computational costs. Recent work (An et al. 2023; Wang et al. 2023; Wang, Li, and Vasconcelos 2021) has highlighted drawbacks in NST methods that rely on pre-trained CNNs, e.g., VGG-19 (Simonyan and Zisserman 2014), for feature extraction from the reference image. Wang et al. (Wang, Li, and Vasconcelos 2021) have enhanced non-VGG architectures’ robustness via activation smoothing in stylization loss. An et al. (An et al. 2023) explore alternative architectures, such as GoogLeNet (Szegedy et al. 2015), yet they lack specificity for NST, yielding unsatisfactory stylization outcomes and high memory use. Instead, our objective is productive mobile NST that incorporates aesthetic features.
@@ 3 Method
As depicted in Figure 2, the AesFA architecture comprises three primary components: a content encoder $E_{c}$, an aesthetic feature encoder $E_{aes}$ in conjunction with kernel-prediction networks $\mathcal{K}=\{K_{1},K_{2},K_{3}\}$, and a generator $G$. Specifically, the input contents are encoded via a content encoder $E_{c}$ and subsequently decomposed into two feature maps containing distinct frequency components. Meanwhile, the style images are processed through the aesthetic feature encoder $E_{aes}$ to encapsulate higher-level aesthetic feature information. Employing the aesthetic feature descriptor $\mathcal{W}$, which is encoded by the aesthetic feature encoder, the kernel-prediction networks $\mathcal{K}$ predict aesthetic feature-aware convolutional kernels and biases for each respective spatial resolution. These predictions then are integrated into the generator alongside the decomposed content latent features. Within the generator, the content features merge with the predicted aesthetic feature-aware convolutional kernels and biases for each corresponding frequency using AdaOct. In the generator’s terminal layer, the synthesized high- and low-frequency images are amalgamated to produce a single style-transferred output. To summarize, the overarching pipeline proceeds as follows:
@@
Octave Convolution. A pivotal aspect of the OctConv operator is its capacity to factorize mixed feature maps by their frequencies while concurrently facilitating efficient communication between high- and low-frequency components. Low-frequency feature maps in OctConv have their spatial resolution reduced by one octave (Lindeberg 2013), where an octave is a spatial dimension divided by a power of two. In this study, a value of 2 was chosen for simplicity, but it can be an arbitrary value. Reducing the spatial resolution in the low-frequency branch expands the receptive field, allowing smooth, low-frequency maps to be stored in a low-resolution tensor. OctConv’s multi-frequency representation effectively separates input feature maps by frequency, reducing spatial redundancy. Given input and output of OctConv as $\mathit{X=\{X_{H},X_{L}\}}$ and $\mathit{Y=\{Y_{H},Y_{L}\}}$, the forward pass is defined as:
@@
Figure 3: Detailed design of the Adaptive Octave convolutions (AdaOct) used in AesFA.
Content and Aesthetic Feature Encoders. Both proposed encoders improve upon MobileNet (Howard et al. 2017) by replacing all convolutions with OctConv to factorize feature maps by their frequency, thereby reducing network redundancy while maintaining simplicity and efficiency. Active information exchange occurs between frequency components during training. Spatial reduction in the low-frequency branch expands the receptive field, capturing more contextual information from distant locations and improving performance with greater image resolution. In contrast to the original OctConv, the upsampling order is adjusted to address checkerboard artifacts.
@@
Kernel-Prediction Networks. To effectively apply the aesthetic feature descriptor to content latent features, we present an approach using kernel-prediction networks similar to those of AdaConv. These networks predict aesthetic feature-aware kernels and biases in a depthwise-separable manner, corresponding to frequency and spatial resolution. The aesthetic feature-aware kernels and biases comprise depthwise convolution components, pointwise convolution components, and per-channel biases. This approach diverges from the original kernel-prediction network utilized in AdaConv by predicting aesthetic feature-aware kernels and biases from both high- and low-frequency aesthetic feature descriptors.
@@
A singular perceptual loss is insufficient for extracting and expressing intricate aesthetic-style representations (refer to Figure  8). To address this limitation, we adopt and improve the contrastive learning approach from MicroAST (Wang et al. 2023). Our enhanced loss, termed Aesthetic Feature Contrastive Loss ($\mathcal{L}_{Aes}$), follows the contrastive learning principle of maintaining proximity between data and their corresponding “positive” samples while distancing them from other instances deemed as “negatives” in the representation space. Consequently, the selection of “positive” and “negative” samples is pivotal for the success of contrastive learning.
@@
Perceptual Loss. In accordance with previous studies (Gatys, Ecker, and Bethge 2016; Johnson, Alahi, and Fei-Fei 2016), we use the pre-trained VGG-19 model to compute the perceptual loss, which consists of both the content and style losses. However, this work redefines the style loss $\mathcal{L}_{S}$ as in the case of EFDM. Given $I$ representing the stylized output and $y$ denoting the style image, the final perceptual losses are as follows:
@@
Figure 6: Qualitative comparison with various NST algorithms in 512 (first and second row) and 2K (third and fourth row) resolution. First row: tones, second row: texture, third row: brushstrokes, and fourth row: grains. In all aesthetic features, our AesFA method outperforms.
4 Experimental Results
@@
Qualitative Comparisons. As described in Figure 5, AesFA qualitatively outperforms eight state-of-the-art NST techniques in terms of aesthetics while maintaining the essential content semantics. AesFA excels in the transfer of unique local aesthetic structural elements from the style image to the content image at all spatial resolutions. Notably, Figure 6 demonstrates that our method can faithfully show aesthetic feature-aware style transfer in terms of tones (first row), texture (second row), brushstrokes (third row), and grains (fourth row). Figure 7 also shows that AesFA excels in transferring the local structure of the style image to the content image in ultra-high resolution (e.g., 4K). MicroAST, for example, suffers from poor aesthetic stylizations and low image quality (blue box in Figure 7). In contrast, AesFA achieves promising outputs with high image quality (red box in Figure 7). Additional results are demonstrated in the supplementary materials.
Figure 7: Ultra-high resolution (4K; 4096$\times$4096) comparison. The top-left images are content and style images, and the top right displays our synthesized image. The bottom-left image is the magnified image of our result, and the image on the right is the magnified result from the current state-of-the-art model. Our model outperforms in terms of aesthetic features (e.g., brushstrokes; texture; tones). Magnify an image to see the details.
Resolution (px) Method Style Loss($\downarrow$) LPIPS($\downarrow$) SSIM($\uparrow$) Time (sec, $\downarrow$) Pref. (%$\uparrow$)
@@
Table 1: Quantitative comparison with various state-of-the-art NST algorithms. “N/A” and “OOM” denote “Not applicable at this resolution” and “Out of GPU memory,” respectively.
@@
User Study. Evaluating the outcomes of stylization is a highly subjective matter. Hence, we have conducted a user study for the nine approaches. We randomly show each participant 20 ballots (4 ballots for each resolution) containing the content, style, and nine outputs. For each ballot, participants were given unlimited time to select their favorite output in terms of aesthetically pleasing stylization and content preservation. We collected 1580 valid votes from 79 subjects. The preference percentage of each method for each resolution is included in the last column of Table 1. The user study results demonstrate that our stylized images are more appealing than or at least comparable to the competitors.
@@
Figure 9: The style blended results generated by AesFA with different style images. “Low” and “High” represent the low- and high-frequency styles, respectively. The main idea of style blending is to mix the style among different style images. Because AesFA effectively decomposes the input image into two different frequency bands, one can choose which image attributes to blend for frequency analysis.
5 Conclusion

### 22 Feb 2024 v3

@@ Introduction
Figure 1: Top: The Starry Night by Vincent Van Gogh. The styles have a strong correlation with spatial information, as evidenced by the presence of whirling patterns and expressionistic yellow stars in the “sky”. Bottom: Compared with other NST methods, our method can faithfully transfer styles while ensuring the spatial information.
@@
Related Work
NST emerged with Gatys et al. (Gatys, Ecker, and Bethge 2016), but its optimization is computationally intensive. To deal with this issue, Johnson et al. (Johnson, Alahi, and Fei-Fei 2016) introduced perceptual losses for real-time processing. Subsequent work (Gatys et al. 2017; Ghiasi et al. 2017; Chen and Schmidt 2016; Ulyanov, Vedaldi, and Lempitsky 2017; Dumoulin, Shlens, and Kudlur 2016; Ulyanov et al. 2016) improved NST without sacrificing speed. However, all were limited to specific styles. Huang et al. (Huang and Belongie 2017) proposed Adaptive Instance Normalization (AdaIN) for arbitrary style transfer, which has been extended (Sheng et al. 2018; Kotovenko et al. 2019; Jing et al. 2020; Shen, Yan, and Zeng 2018; Li et al. 2017; lin2021drafting; yoo2019photorealistic) for successful style transfer onto content images. Chandran et al. (Chandran et al. 2021) improved AdaIN with AdaConv for structure-aware style transfer. AdaConv simultaneously adapts statistical and structural styles. However, AdaConv’s convolution kernels and biases incur high computational costs. Recent work (An et al. 2023; Wang et al. 2023; Wang, Li, and Vasconcelos 2021) has highlighted drawbacks in NST methods that rely on pre-trained CNNs, e.g., VGG-19 (Simonyan and Zisserman 2014), for feature extraction from the reference image. Wang et al. (Wang, Li, and Vasconcelos 2021) have enhanced non-VGG architectures’ robustness via activation smoothing in stylization loss. An et al. (An et al. 2023) explore alternative architectures, such as GoogLeNet (Szegedy et al. 2015), yet they lack specificity for NST, yielding unsatisfactory stylization outcomes and high memory use. Instead, our objective is productive mobile NST that incorporates aesthetic features.
@@
As depicted in Figure 2, the AesFA architecture comprises three primary components: a content encoder $E_{c}$, an aesthetic feature encoder $E_{aes}$ in conjunction with kernel-prediction networks $\mathcal{K}$, and a generator $G$. Specifically, the input contents are encoded via a content encoder $E_{c}$ and subsequently decomposed into two feature maps containing distinct frequency components. Meanwhile, the style images are processed through the aesthetic feature encoder $E_{aes}$ to encapsulate higher-level aesthetic feature information. Employing the aesthetic feature descriptor $\mathcal{W}$, which is encoded by the aesthetic feature encoder, the kernel-prediction networks $\mathcal{K}$ predict aesthetic feature-aware convolutional kernels and biases for each respective spatial resolution. These predictions then are integrated into the generator alongside the decomposed content latent features. Within the generator, the content features merge with the predicted aesthetic feature-aware convolutional kernels and biases for each corresponding frequency using AdaOct. In the terminal layer, the synthesized high- and low-frequency images are amalgamated to produce a single style-transferred output. To summarize, the overarching pipeline proceeds as follows:
@@
Octave Convolution. A pivotal aspect of the OctConv operator is its capacity to factorize mixed feature maps by their frequencies while concurrently facilitating efficient communication between high- and low-frequency components. Low-frequency feature maps in OctConv have their spatial resolution reduced by one octave (Lindeberg 2013), where an octave is a spatial dimension divided by a power of two. In this study, a value of 2 was chosen for simplicity. Given input and output of OctConv as $\mathit{X=\{X_{H},X_{L}\}}$ and $\mathit{Y=\{Y_{H},Y_{L}\}}$, the forward pass is defined as:
@@
Figure 3: The detailed design of the Adaptive Octave Convolutions (AdaOct) used in AesFA.
Content and Aesthetic Feature Encoders. Both proposed encoders improve upon MobileNet (Howard et al. 2017) by replacing all convolutions with OctConv to factorize feature maps by their frequency, thereby reducing network redundancy while maintaining simplicity and efficiency. Spatial reduction in the low-frequency branch expands the receptive field, capturing more contextual information from distant locations and improving performance with greater image resolution. In contrast to the original OctConv, the upsampling order is adjusted to address checkerboard artifacts.
@@
Kernel-Prediction Networks. To effectively apply the aesthetic feature descriptor to content features, we present an approach using kernel-prediction networks similar to those of AdaConv. These networks predict aesthetic feature-aware kernels and biases in a depthwise-separable manner, corresponding to frequency and spatial resolution. The aesthetic feature-aware kernels and biases comprise depthwise convolution components, pointwise convolution components, and per-channel biases. This approach diverges from the original kernel-prediction network utilized in AdaConv by predicting aesthetic feature-aware kernels and biases from both high- and low-frequency aesthetic feature descriptors.
@@
A singular perceptual loss is insufficient for extracting and expressing intricate aesthetic-style representations (refer to Figure 8). To address this limitation, we adopt and improve the contrastive learning approach from MicroAST (Wang et al. 2023). Our enhanced loss, termed Aesthetic Feature Contrastive Loss ($\mathcal{L}_{Aes}$), follows the contrastive learning principle of maintaining proximity between data and their corresponding “positive” samples while distancing them from other instances deemed as “negatives” in the representation space. Consequently, the selection of “positive” and “negative” samples is pivotal for the success of contrastive learning.
@@
Perceptual Loss. In accordance with previous studies (Gatys, Ecker, and Bethge 2016; Johnson, Alahi, and Fei-Fei 2016), we use the pre-trained VGG-19 model to compute the perceptual loss, which consists of both the content and style losses. However, this work redefines the style loss $\mathcal{L}_{S}$ as in the case of EFDM. Given $I$ representing the stylized output and $y$ denoting the reference image, the final perceptual losses are as follows:
@@
Figure 6: Qualitative comparison with various NST algorithms in 512 (first and second row) and 2K (2048$\times$2048; third and fourth row) resolution. First row: tones, second row: texture, third row: brushstrokes, and fourth row: grains. In all aesthetic features, our AesFA method outperforms.
Experimental Results
@@
Qualitative Comparisons. As described in Figure 5, AesFA qualitatively outperforms eight state-of-the-art NST techniques in terms of aesthetics while maintaining the essential content semantics. AesFA excels in the transfer of unique local aesthetic structural elements from the style image to the content image at all spatial resolutions. Notably, Figure 6 demonstrates that our method can faithfully show aesthetic feature-aware style transfer in terms of tones (first row), texture (second row), brushstrokes (third row), and grains (fourth row). Figure 7 also shows that AesFA excels in transferring the local structure of the style image to the content image in ultra-high resolution (e.g., 4K). MicroAST, for example, suffers from poor aesthetic stylizations and low image quality (blue box in Figure 7). In contrast, AesFA achieves promising outputs with higher image quality (red box in Figure 7). Additional results are demonstrated in the supplementary materials.
Figure 7: Ultra-high resolution (4K; 4096$\times$4096) comparison. The top-left images are content and style images, and the top right displays our synthesized image. The bottom-left image is the magnified image of our result, and the image on the right is the magnified result from the current state-of-the-art model. Our model outperforms in terms of aesthetic features (e.g., brushstrokes; texture; tones). Zoom in for details.
Resolution Method Style Loss ($\downarrow$) LPIPS ($\downarrow$) SSIM ($\uparrow$) Time (sec, $\downarrow$) Pref. (%, $\uparrow$)
@@
Table 1: Quantitative comparison with various state-of-the-art NST algorithms. “N/A” means “Not applicable at this resolution” and “OOM” stands for “Out of GPU memory”.
@@
User Study. Evaluating the outcomes of stylization is a highly subjective matter. Hence, we have conducted a user study for the nine approaches. We randomly show each participant 20 ballots (4 ballots for each resolution) containing the content, style, and nine outputs. For each ballot, participants were given unlimited time to select their favorite output in terms of aesthetically pleasing stylization and content preservation. We collected 1,580 valid votes from 79 subjects. The preference percentage of each method for each resolution is included in the last column of Table 1. The user study results demonstrate that our stylized images are more appealing than or at least comparable to the competitors.
@@
Figure 9: The style blending results generated by AesFA with different style images. Zoom in for details.
Figure 9 shows the style blending, i.e., using the low- and high-frequency style information from different style images. Sub-figures (a)-(d) show different combinations of origins for low- and high-frequency style information. For instance, “(b) Low-S1 / High-S2” indicates that we use the low-frequency style information from image “S1” and the high-frequency style information from image “S2”.
Conclusion